Question 1

Solve the following equation

a) 2e^2x−1 + 5e^x2= 0

b) 2^4x + 2^2x−1 > 2

[8,6]

Question 2

Let f : (−∞, 2] → R be given by

f(x) = √2 − x

a) Show that f is injective.

b) Determine im(f).

c) Find a left inverse g : R → (−∞, 2] of f.

[7,6,5]


Question 3

Let f : Z → Z be given by

f(z) = (2z − 5, if z ≥ 0;

(z + 5, if z < 0.

a) Is f an injective function?

b) Let u ∈ Z, u ≤ 5. show that u ∈ im(f).

c) Let v ∈ Z, v > 5, Show that v ∈ im(f) if and only if v + 5 is even .

[5,6,6]



1
Expert's answer
2021-10-18T13:05:02-0400

Solution.

Question 1

a)

equation "2e^{2x-1}+5e^{x^2}=0" has not real roots because "a^{f(x)}>0, a>0, a\neq1" for all x.

b)

"2^{4x}+2^{2x-1}>2"

"2\u20222^{4x}+2^{2x}-2\u20222>0"

Let be "t=2^{2x}," then

"2t^{2}+t-4>0"

From here "t<\frac{-1-\sqrt{33}}{4}," and "t>\frac{-1+\sqrt{33}}{4}."

"2^{2x}<\frac{-1-\sqrt{33}}{4}" has not real solutions.

"2^{2x}>\frac{-1+\sqrt{33}}{4},"

"2x>\log_2{\frac{-1+\sqrt{33}}{4}},"

"x>\frac{1}{2}\log_2\frac{-1+\sqrt{33}}{4}."

Question 2

"f:(-\infty,2]\to R, f(x)=\sqrt{2-x}"

a) each x from "(-\infty,2)" has y to correspond, for a example, f(0)="\sqrt{2}," f(-2)=2, f(-7)=3.



So, f(x) is injective function.

b) as "x\in (-\infty,2]" f(x) is an increasing function in the domain

"x\to -\infty, f(x)\to +\infty"

So, Im(f)="+\infty."

c) find a left inverse of f:

"f^2=2-x,\newline
x=2-f^2,\newline
g(x)=-x^2+2, x\in [0,+\infty)."



Question 3

"f:Z\to Z, f(z)=2z-5, if z\geq0; z+5, if z<0."

a) "f(z)=0, z=2.5, \text{or } z=-5."

So, f(z) is not injective.

b) such as "z\leq 5," we have

"z\in (-\infty,5]\implies u\leq 5 \in" im(f).

c) "v\in Z,v>5."

Let be "v=7," 7+5=12, 12 is even.